May I please have some help? A soccer field is 336yd^2. The width is 5yd longer than its length. What is the length and width of the soccer field. Thanks for your help.
Answer by mszlmb(115)

Given soccer fields are rectangular:
2 variables, the width, and the length of the soccer field.
The width is 5 longer than the length.
The width times the length is 336.
The unit is yards.
W=L+5
WL=336
substitute for W in "WL=336".
(L+5)L=336 distribute
L2+5L=336 subtract 336 from both sides
L2+5L-336=0 Factor (or use quadradic formula)
2 numbers which multiply to -336 and add to 5, so 1 number must be negative.
336=112*3=56*6=28*12=14*24=7*48, but you'll notice none of those have a difference of 5..let's try 2's
336=168*2=84*4=42*8=21*16 aha! difference of 5. so we alter the equation:
L2+5L-336=L2-16L+21L-336
Now we factor:
L2-16L+21L-336 take out like terms:
L(L-16)+21(L-16) put like terms 2gether
(L+21)(L-16)=0 so 1 of thoz = 0
L=-21, or it equals 16. A side is hardly negative; but definitely not for a football field, therefore L=16.
W=L+5 so W=16+5, or 21.
W=21yards
L=15yards
..I'm beginning to think there was a simpler way of figuring this out, I cannot put my finger on it.